OpenStudy (brucebaner):

Choose all answers that are correct. 1. (1 pt) Relationship B has a greater rate than Relationship A. The graph represents Relationship A. Which equations could represent Relationship B? Choose all answers that are correct.

OpenStudy (brucebaner):

A: y=3/4x B: y=0.6x C: y=2/3x D: y=1/4x

OpenStudy (brucebaner):

@chosenmatt

OpenStudy (brucebaner):

can you plz help me @chosenmatt

OpenStudy (hyper):

Firstly, calculate the slope of the graph .... The formula for the slope of the graph is: m = slope = \[\frac{ (y_{2} - y _{1}) }{ (x _{2} - x _{1}) }\]

OpenStudy (hyper):

Have you calculated the slope for the graph which represents relationship A?